OpenStudy (anonymous):

Candice deposited $3,587.39 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $5,000? 8 years, 0 months 16 years, 7 month 8 years, 1 month 9 years, 10 months

hero (hero):

Semi-Annual Compounding Formula \(A = P\left(1 + \frac{r}{2}\right)^{2t}\)

hero (hero):

A = 5000 P = 3587.39 r = .042 \(5000 = 3587.39\left(1 + \frac{.042}{2}\right)^{2t}\) Solve for \(t\)

OpenStudy (anonymous):

I got 0

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

\[\Large 5000 = 3587.39\left(1 + \frac{.042}{2}\right)^{2t}\] first divide both sides by 3587.39 do you know how to take logs of both sides?

OpenStudy (agent0smith):

After dividing both sides by 3587.39 \[\Large 1.39377= \left(1 + \frac{.042}{2}\right)^{2t}\]now you have to take logs of both sides